OpenStudy (anonymous):

simplify first. then solve using logs. 1322.17/180000=(.03897/12)/(1-(1+(.03897/12))^(-12t))

OpenStudy (anonymous):

\[t\to \frac{\text{Log}\left[\frac{132217}{73762}\right]}{12 \text{Log}\left[\frac{401299}{400000}\right]}= 14.999947 \]